# Advantage Shaping as Surrogate Reward Maximization:

### v1

@@ Abstract
This note reconciles two seemingly distinct approaches to policy gradient optimization for the Pass@K objective in reinforcement learning with verifiable rewards: (1) direct REINFORCE-style methods, and (2) advantage-shaping techniques that directly modify GRPO. We show that these are two sides of the same coin. By reverse-engineering existing advantage-shaping algorithms, we reveal that they implicitly optimize surrogate rewards. We specifically interpret practical “hard-example up-weighting” modifications to GRPO as reward-level regularization. Conversely, starting from surrogate reward objectives, we provide a simple recipe for deriving both existing and new advantage-shaping methods. This perspective provides a lens for RLVR policy gradient optimization beyond our original motivation of Pass@K.
@@ 1 Introduction
When evaluating large language models on math and coding tasks, standard practice is to generate $K$ independent solutions and check whether any succeeds; the so-called Pass@K metric [Kul+19, Che+21]. Yet most policy gradient methods, from REINFORCE [Wil92] to more recent variants like RLOO [KHW19, Ahm+24] and GRPO [Sha+24], optimize for single-attempt reward creating a mismatch. Should we instead train what we test?
Recent work has approached this question from two seemingly distinct angles. [Mah+25, WK25] derive policy gradients that directly maximize Pass@K reward, yielding REINFORCE-style updates with reweighted advantages. Meanwhile, [Che+25] propose modifying GRPO through advantage shaping, that is, directly adjusting advantage scores to account for the Pass@K objective. Both improve Pass@K performance, but their relationship remains unclear. Are these fundamentally different techniques, or two views of the same underlying principle?
This note unifies advantage shaping and direct optimization, showing they are two sides of the same coin. As a byproduct, we establish reward regularization as a design mechanism for policy gradient optimization with verifiable rewards, providing a recipe for algorithm interpretation via reverse-engineering and development via forward-engineering.
@@ 1.1 Problem Setup: Reinforcement Learning with Verifiable Rewards
We study reinforcement learning with verifiable rewards (RLVR). We are given a distribution $\mathcal{P}$ over problem-answer pairs $(x,a)$, where $x$ might be a math problem and $a$ its final numeric answer. The goal is to train a large language model (LLM) to maximize the reward of finding correct answers to problems from $\mathcal{P}$.
We abstract the LLM as a conditional distribution $\pi_{\theta}(\text{response}|\text{prompt})$ parameterized by $\theta$. Given a prompt $x$, the model generates a response $y$ (a sequence of tokens) by autoregressive sampling. We seek parameters $\theta$ that maximize the expected reward:
@@
1.1.1 The 0/1 Reward Baseline
The standard reward function is the binary 0/1 reward:
@@ 1.1.1 The 0/1 Reward Baseline
A popular algorithm for maximizing expected 0/1 reward is REINFORCE Leave-One-Out (RLOO) [KHW19, Ahm+24], which extends the classical REINFORCE algorithm [Wil92] using multiple generated responses. At each iteration $t$, and for each pair $(x,a)$ from a training set $\mathcal{D}$, RLOO updates the model parameters as
@@
Despite its simplicity, RLOO is a strong baseline [Ahm+24]. Its structure also underlies recent variants like GRPO [Sha+24]. As noted by [Den+25], in full online mode (when clipping can be ignored), GRPO uses the same form as (RLOO) with normalized advantages $A_{i}^{\text{GRPO}}=A_{i}^{\text{RLOO}}/\sqrt{\hat{\rho}(1-\hat{\rho})}$, where $\hat{\rho}$ is the empirical 0/1 reward for example $(x,a)$.
@@ 1.1.2 The Pass@K Reward
The 0/1 reward, however, misaligns with practical LLM usage. Given a problem $x$, users typically prompt the model multiple times (say, $K$ times) and examine the responses $y_{1},\ldots,y_{K}$ to find one that produces the desired answer. Accordingly, LLMs are often evaluated using the Pass@K metric [Che+21], which measures whether at least one solution is correct among $K$ generated responses:
@@
Since test-time performance is measured via Pass@K, a natural question arises: why not train the model to directly maximize Pass@K reward rather than 0/1 reward? That is, the learning goals becomes maximizing $\mathbb{E}_{(x,a)\sim\mathcal{P}}[\rho_{K,\theta}(x,a)]\,$, where we define the per-example expected Pass@K reward as
@@ 1.2 Contributions
C-1. Direct Pass@K optimization. Building on [Mah+25], we show that the exact Pass@K policy gradient of $\rho_{K,\theta}=\rho_{K,\theta}(x,a)$ equals the 0/1 gradient reweighted by per-example Fail@(K-1) probabilities. This yields unbiased Monte Carlo estimators REINFORCEK/RLOOK using leave-one-out weights, and biased variants GRPOK. All methods naturally upweight rare successes and can be cast as asymmetric reweightings of their 0/1-reward analogues.
C-2. Advantage shaping == surrogate reward maximization. We show that the GRPO-based approach of [Che+25] can be reverse-engineered to reveal that it asymptotically (in the number of generated responses) maximizes a surrogate reward $\frac{2}{K}\arcsin(\sqrt{\rho_{K,\theta}})$. We also provide forward engineering that reproduces advantage-shaping heuristics in three steps:
@@
    Differentiate the surrogate reward using the log-derivative trick;
@@
    Substitute all population quantities with their empirical analogues;
  3. (iii)
    Replace the population reward gradient with its RLOO proxy (Eq. (1)).
As a corollary, when K=1K{=}1, vanilla GRPO is (up to min-clip operations) effectively RLOO applied to the surrogate reward $2\arcsin(\sqrt{\rho_{\theta}})$, a variance-stabilizing transformation of the 0/1 reward.
C-3. Reward-level regularization. We interpret commonly used heuristics that downweight “easy” examples (that is, those already solved with high probability) and upweight “hard” ones [Che+25, Tea+25]. We show that multiplying GRPO gradient contributions by factors such as $1-\widehat{\rho}$ is effectively equivalent to optimizing a regularized surrogate reward $\arcsin(\sqrt{\rho_{\theta}})+\sqrt{\rho_{\theta}(1-\rho_{\theta})}$. This can be interpreted as a special instance of regularized surrogate reward maximization:
@@
where the stochastic training objective over examples $(x,a)\sim\mathcal{D}$ consists of a data-fitting term $F$, that a strong baseline like GRPO would directly optimize, and, an additive regularizer $\Omega$ (balanced via $\lambda\geq 0$). Unlike typical policy-level or parameter-level regularization, this operates simply at the reward level. We demonstrate the regularization perspective by deriving entropy-regularized variants using the three-step recipe outlined in Contribution C-2, yielding new tunable advantage-shaping algorithms.
@@
For a given example $(x,a)$, let $y_{1},\ldots,y_{N}$ denote $N$ responses generated IID from the model $\pi_{\theta}(\cdot|x)$. Let $r_{i}:=r_{0/1}(y_{i},a)$ denote the 0/1 reward of response $i$, and let $\widehat{\rho}:=\frac{1}{N}\sum_{i\in[N]}r_{i}$ denote the empirical 0/1 reward. Let $N^{+}:=\sum_{i\in[N]}r_{i}$ and $N^{-}:=N-N^{+}$ denote the number of correct and wrong responses, respectively. We use the shorthand $\nabla_{i}:=\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ for the log-probability gradient of individual response $y_{i}$, and for the average log-probability gradient over correct/wrong responses:
@@ 2 Warm-up: Optimizing the 0/1 Reward
REINFORCE-style algorithms provide different approximations of the per-example expected gradient $G_{0/1}(\theta;(x,a))=\nabla_{\theta}\rho_{\theta}(x,a)$.
@@
The classical REINFORCE algorithm [Wil92] approximates the expectation over model outputs in (4) using the empirical average over $N$ sampled responses. This yields the per-example gradient update:
@@
REINFORCE with Leave-One-Out (RLOO) [KHW19, Ahm+24] reduces variance by replacing rewards $r_{i}$ with advantage scores that subtract a leave-one-out baseline:
@@
Group Relative Policy Optimization (GRPO) [Sha+24] normalizes advantages by their empirical standard deviation:
@@
Ignoring the clipping operation, the GRPO gradient is effectively [Den+25]:
@@
3 Pass@K via Direct Differentiation
@@
Following [Mah+25], we apply the log-derivative trick:
@@
$G_{\text{\tiny Pass@K}}(\theta;(x,a))$ $:=K\cdot\underbrace{\left(1-\rho_{K-1,\theta}(x,a)\right)}_{\text{Fail@(K-1)}}\,\cdot\,\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{0/1}(y,a)\nabla_{\theta}\log\pi_{\theta}(y|x)\right]\,.$ (11)
This is a reweighting of the respective gradient for the 0/1 reward. For $K=1$, Pass@1 equals the 0/1 reward and the gradient updates are indeed equivalent.
@@
To arrive at an unbiased estimator, we first require an unbiased estimator for the Fail@(K-1) term in Eq. (11). The following is an unbiased estimator of $\rho_{K,\theta}(x,a)$ [Che+21]:
@@
Analogously, we obtain a GRPO update for Pass@K optimization by using the GRPO advantage scores from Eq. (8):
@@ 4 Pass@K by Advantage Shaping
We now review the GRPO variant proposed by [Che+25], which we denote $\widetilde{\text{GRPO}}_{K}$, and which directly shapes GRPO’s advantages to favor Pass@K improvements. We refer readers to their nicely-written paper for details of the design rationale. Here, we begin from their proposed formulas for the modified GRPO advantages, which we denote $\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}$. [Che+25, App. B3 ] propose the following scores:
@@
As noted in [Den+25a], this can be simplified: Recall from Eq. (14) that $1-\widehat{\rho}_{K}=(1-\widehat{\rho})\cdot\widehat{f}_{K-1}^{-}$, which reveals $\widetilde{A}_{\text{\tiny Pass@K}}^{-}=-\frac{\widehat{\rho}}{1-\widehat{\rho}}\cdot\widetilde{A}_{\text{\tiny Pass@K}}^{+}$. Thus, in terms of vanilla GRPO advantage scores:
@@ 5.1 Reverse-Engineering: Surrogate Reward from Advantage Shaping
Recall $\widehat{\rho}:=\widehat{\rho}_{\theta}(x,a)$ denotes the per-example empirical 0/1 reward and $\widehat{\rho}_{K}:=\widehat{\rho}_{K,\theta}(x,a)$ denotes the per-example empirical estimate of the Pass@K reward (Eq. (12)). The method of [Che+25], $\widetilde{\text{GRPO}}_{K}$, rescales the vanilla GRPO empirical gradient yielding updates as shown in Eq. (21). For sufficiently large sample size $N\gg K$, this corresponds to direct maximization of the per-example surrogate reward
@@
Thus, the advantage-shaping approach of [Che+25] implicitly optimizes a smooth, differentiable transformation of the Pass@K reward, providing further justification for the validity of the different approach by which the authors arrive at it. Note that the surrogate reward is maximized when the Pass@K reward equals 1, and is actually strictly monotone, so both the original and surrogate objectives share the same optimal solution, though the arcsin surrogate reward leads to a different optimization path.
@@
The arcsin transform that emerges is noteworthy: it is a variance-stabilizing transformation (VST) for the binomial distribution. Specifically, for $\mathcal{X}\sim\text{Bin}(M,p)$, the arcsin transform satisfies $\text{Var}(\sqrt{M}\cdot\arcsin(\sqrt{\mathcal{X}/M}))\approx 1/4$ independent of $p$ [Ans48]. This connection is not coincidental: [Che+25, Sec. 2.2 ] arrived at their method through arguments using batched sampling, where $N$ responses are partitioned into $N/K$ groups of size $K$. In this setting, the number of successful groups follows a binomial distribution with $M=N/K$ trials, and the arcsin transform would stabilize the variance of this estimate. Formalizing the connection between VSTs and optimization stability when applying empirical gradients to transformed rewards is open direction for future work.

### v2

@@ Abstract
We reconcile two seemingly distinct approaches to policy gradient optimization for the Pass@K objective in reinforcement learning with verifiable rewards: (1) direct REINFORCE-style methods, and (2) advantage-shaping techniques that directly modify GRPO. We show that these are two sides of the same coin. By reverse-engineering existing advantage-shaping algorithms, we reveal that they implicitly optimize surrogate rewards. We specifically interpret practical “hard-example upweighting” modifications to GRPO as reward-level regularization. Conversely, starting from surrogate reward objectives, we provide a simple recipe for deriving both existing and new advantage-shaping methods. This perspective provides a lens for RLVR policy gradient optimization beyond our original motivation of Pass@K.
@@ 1 Introduction
When evaluating large language models on math and coding tasks, standard practice is to generate $K$ independent solutions and check whether any succeeds; the so-called Pass@K metric [kulal2019spoc, codex]. Yet most policy gradient methods, from REINFORCE [williams1992simple] to more recent variants like RLOO [kool2019buy, ahmadian2024back] and GRPO [GRPO_original], optimize for single-attempt reward creating a mismatch. Should we instead train what we test?
Recent work has approached this question from two seemingly distinct angles. [sadegh, passk_unbiased_gradient] derive policy gradients that directly maximize Pass@K reward, yielding REINFORCE-style updates with reweighted advantages. Meanwhile, [passk_bydance] propose modifying GRPO through advantage shaping, that is, directly adjusting advantage scores to account for the Pass@K objective. Both improve Pass@K performance, but their relationship remains unclear. Are these fundamentally different techniques, or two views of the same underlying principle?
This work unifies these seemingly distinct perspectives, showing that advantage shaping and direct optimization are two sides of the same coin. The key is recognizing that different policy gradient methods can be understood as optimizing different surrogate rewards; see Section 1.2. As a byproduct, we establish reward-level regularization as a design mechanism for policy gradient optimization with verifiable rewards, providing a recipe for algorithm interpretation via reverse-engineering and development via forward-engineering.
@@ 1.1 Problem Setup: Reinforcement Learning with Verifiable Rewards
We study reinforcement learning with verifiable rewards (RLVR) [RLVR_term, GRPO_original, r1], a framework for training Large Language Models (LLMs) on tasks with objectively verifiable outcomes (see [RLLLMsurvey2025] for a recent survey). Concretely, we are given a distribution $\mathcal{P}$ over problem-answer pairs $(x,a)$, where $x$ might be a math problem and $a$ its final numeric answer. The goal is to train an LLM to maximize the reward of finding correct answers to problems from $\mathcal{P}$. We abstract the LLM as a conditional distribution $\pi_{\theta}(\text{response}|\text{prompt})$ parameterized by $\theta$. Given a prompt $x$, the model generates a response $y$ (a sequence of tokens) by autoregressive sampling. We seek parameters $\theta$ that maximize the expected reward:
@@
1.1.1 The 0/1 Reward
The most popular reward function in RLVR is the binary 0/1 reward:
@@ 1.1.1 The 0/1 Reward
A popular algorithm for maximizing expected 0/1 reward is REINFORCE Leave-One-Out (RLOO) [kool2019buy, ahmadian2024back], which extends the naive REINFORCE algorithm [williams1992simple] using multiple generated responses to form a baseline. At each iteration $t$, and for each pair $(x,a)$ from a training set $\mathcal{D}$, RLOO updates the model parameters as
@@
Despite its simplicity, RLOO is a strong baseline [ahmadian2024back]. Its structure also underlies recent popular variants like GRPO [GRPO_original]. In full online mode (when clipping can be ignored), GRPO uses the same form as (RLOO) with normalized advantages $A_{i}^{\text{GRPO}}=A_{i}^{\text{RLOO}}/\sqrt{\hat{\rho}(1-\hat{\rho})}$, where $\hat{\rho}$ is the empirical 0/1 reward for example $(x,a)$.
@@ 1.1.2 The Pass@K Reward
The 0/1 reward, however, misaligns with practical LLM usage. Given a problem $x$, users typically prompt the model multiple times (say, $K$ times) and examine the responses $y_{1},\ldots,y_{K}$ to find one that produces the desired answer. Accordingly, LLMs are often evaluated using the Pass@K metric [codex], which measures whether at least one solution is correct among $K$ generated responses:
@@
Since test-time performance is measured via Pass@K, we may train the model to directly maximize Pass@K reward rather than 0/1 reward by maximizing $\mathbb{E}_{(x,a)\sim\mathcal{P}}[\rho_{K,\theta}(x,a)]\,$, where we define the per-example expected Pass@K reward as
@@ 1.2 Contributions
C-1. Direct Pass@K optimization. Building on [sadegh], we show that the exact Pass@K policy gradient equals the 0/1 gradient reweighted by per-example Fail@(K-1) probabilities. This yields Monte Carlo unbiased estimators REINFORCEK/RLOOK and a biased variant GRPOK, all expressible as asymmetric reweightings of their 0/1-reward counterparts.
C-2. Advantage shaping == surrogate reward maximization. We reverse-engineer the advantage-shaping algorithm of [passk_bydance] and find that it asymptotically (in the number of generated responses) maximizes a surrogate reward $\frac{2}{K}\arcsin(\sqrt{\rho_{K,\theta}})$. Conversely, we provide a forward-engineering recipe: starting from any differentiable surrogate reward $F(\rho_{\theta})$, advantage-shaping heuristics can be systematically derived as follows:
@@
    Differentiate the surrogate reward: $\nabla_{\theta}F(\rho_{\theta})=F^{\prime}(\rho_{\theta})\cdot\nabla_{\theta}\rho_{\theta}$
@@
    Substitute all population quantities with their empirical analogues:
    1. (a)
[…]
    2. (b)
      Replace the population reward gradient $\nabla_{\theta}\rho_{\theta}$ with its RLOO proxy $\widehat{G}_{\text{RLOO}}(\theta)$ (Eq. (1)).
As a corollary, when K=1K{=}1, vanilla GRPO is (up to clipping operations) effectively RLOO applied to the surrogate reward $F(\rho_{\theta})=2\arcsin(\sqrt{\rho_{\theta}})$, a variance-stabilizing transformation of the 0/1 reward.
C-3. Reward-level regularization. We interpret commonly used heuristics that downweight “easy” examples (those already solved with high probability) and upweight “hard” ones [passk_bydance, team2025kimi]. We show that multiplying GRPO gradient contributions by factors such as $1-\widehat{\rho}$ is effectively equivalent to optimizing a regularized surrogate reward $\arcsin(\sqrt{\rho_{\theta}})+\sqrt{\rho_{\theta}(1-\rho_{\theta})}$. More generally, this can be interpreted as a special instance of regularized surrogate reward maximization:
@@
where the stochastic training objective over examples $(x,a)\sim\mathcal{D}$ consists of two components: a data-fitting term $F$ (e.g., the GRPO surrogate) and an additive regularizer $\Omega$ that encourages maintaining some probability mass on wrong responses, thereby exploring alternative solution paths that may generalize beyond the training set. Unlike typical policy-level or parameter-level regularization, this form of regularization operates simply at the reward level. To demonstrate its use, we instantiate the recipe of Contribution C-2 with entropy regularization, yielding a new tunable advantage-shaping algorithm.
Table 1: Unified view of main algorithms studied in this work. Rows grouped by target evaluation metric. $\rho,\rho_{K}$ are per-example expected 0/1 and Pass@K rewards; $\hat{\rho},\hat{\rho}_{K}$ are their empirical estimators
Algorithm Advantage Scores $(A^{+},A^{-})$ Weighted Empirical Gradient Population Surrogate Reward
[…]
Section 1.2 unifies the main Pass@K policy-gradient algorithms studied in this work. Starting from their advantage scores, we express every method (middle column, ignoring normalization constants and clipping) as $w_{+}\,\widehat{\nabla}_{+}-w_{-}\,\widehat{\nabla}_{-}$, where effective gradient weights $w_{\pm}$ multiply the average empirical gradients over correct/wrong responses. Leveraging this unified representation, we reverse-engineer the corresponding population surrogate reward that each algorithm implicitly optimizes (last column). Conversely, using the forward-engineering recipe from Contribution C-2, we can map a chosen surrogate reward back to advantage scores, establishing a bidirectional relationship:
[…]
@@
For given example $(x,a)$, let $y_{1},\ldots,y_{N}$ denote $N$ responses generated IID from model $\pi_{\theta}(\cdot|x)$. Let $r_{i}:=r_{0/1}(y_{i},a)$ denote the 0/1 reward of response $i$, and $\widehat{\rho}:=\frac{1}{N}\sum_{i\in[N]}r_{i}$ denote the empirical 0/1 reward. Let $N^{+}:=\sum_{i\in[N]}r_{i}$ and $N^{-}:=N-N^{+}$ denote the number of correct and wrong responses. We use the shorthand $\nabla_{i}:=\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ for the log-probability gradient of individual response $y_{i}$, and for the average log-probability gradient over correct/wrong responses:
@@ 2 Warm-up: Optimizing the 0/1 Reward
The REINFORCE-style algorithms reviewed below provide different approximations of the per-example expected gradient $G_{0/1}(\theta;(x,a))=\nabla_{\theta}\rho_{\theta}(x,a)$.
@@
Classical REINFORCE [williams1992simple] approximates the expectation over $y\sim\pi_{\theta}(\cdot|x)$ in (4) using the empirical average over $N$ sampled responses. This yields the per-example gradient:
@@
REINFORCE with Leave-One-Out (RLOO) [kool2019buy, ahmadian2024back] reduces variance by replacing in Eq. (5) rewards $r_{i}$ with advantage scores that subtract a leave-one-out baseline:
@@
Group Relative Policy Optimization (GRPO) [GRPO_original] normalizes advantages by their empirical standard deviation:
@@
Thus, ignoring the clipping operation, which is inactive the fully-online operating mode (see App. A), the GRPO gradient can be expressed conveniently as [deng2025effect]:
@@
3 Pass@K by Direct Differentiation
@@
Following [sadegh], we apply the log-derivative trick:
@@
$G_{\text{\tiny Pass@K}}(\theta;(x,a))$ $:=K\cdot\underbrace{\left(1-\rho_{K-1,\theta}(x,a)\right)}_{\text{Fail@(K-1)}}\,\cdot\,G_{0/1}(\theta;(x,a))\,.$ (11)
This is a reweighting of the respective gradient $G_{0/1}(\theta;(x,a))=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{0/1}(y,a)\nabla_{\theta}\log\pi_{\theta}(y|x)\right]$ for the 0/1 reward. For $K=1$, the gradient updates are indeed equivalent.
@@
To arrive at an unbiased estimator, we require an unbiased estimator for the Fail@(K-1) term in Eq. (11). For this, we leverage the following unbiased estimator of $\rho_{K,\theta}(x,a)$ [codex]:
@@
Analogously, we obtain a GRPO update for Pass@K optimization by using the GRPO advantage scores from Eq. (8) in place of $r_{i}$ in Eq. (13):
@@ 4 Pass@K by Advantage Shaping
We now review the GRPO variant proposed by [passk_bydance], which we denote $\widetilde{\text{GRPO}}_{K}$, and which directly shapes GRPO’s advantages to favor Pass@K improvements. We refer readers to their nicely written paper for details of the design rationale. Here, we begin from their proposed formulas for the modified GRPO advantages, which we denote $\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}$. [passk_bydance, App. B3 ] propose the following scores:
@@
As noted in [thr], this can be simplified: Recall from Eq. (14) that $1-\widehat{\rho}_{K}=(1-\widehat{\rho})\cdot\widehat{f}_{K-1}^{-}$, which reveals $\widetilde{A}_{\text{\tiny Pass@K}}^{-}=-\frac{\widehat{\rho}}{1-\widehat{\rho}}\cdot\widetilde{A}_{\text{\tiny Pass@K}}^{+}$. Thus, in terms of vanilla GRPO advantage scores:
@@ 5.1 Reverse-Engineering: Surrogate Reward from Advantage Shaping
Recall $\widehat{\rho}:=\widehat{\rho}_{\theta}(x,a)$ denotes the per-example empirical 0/1 reward and $\widehat{\rho}_{K}:=\widehat{\rho}_{K,\theta}(x,a)$ denotes the per-example empirical estimate of the Pass@K reward (Eq. (12)). The method of [passk_bydance], $\widetilde{\text{GRPO}}_{K}$, rescales the vanilla GRPO empirical gradient yielding updates as shown in Eq. (21). For sufficiently large sample size $N\gg K$, this corresponds to direct maximization of the per-example surrogate reward
@@
Thus, the advantage-shaping approach of [passk_bydance] implicitly optimizes a smooth, differentiable transformation of the Pass@K reward, providing further justification for the validity of the different approach by which the authors arrive at it. Note that the surrogate reward is maximized when the Pass@K reward equals 1, and is actually strictly monotone, so both the original and surrogate objectives share the same optimal solution, though the arcsin surrogate reward leads to a different optimization path.
@@
The arcsin transform that emerges is noteworthy: it is a variance-stabilizing transformation (VST) for the binomial distribution. Specifically, for $\mathcal{X}\sim\text{Bin}(M,p)$, the arcsin transform satisfies $\text{Var}(\sqrt{M}\cdot\arcsin(\sqrt{\mathcal{X}/M}))\approx 1/4$ independent of $p$ [VST_anscombe1948transformation]. This connection is not coincidental: [passk_bydance, Sec. 2.2 ] arrived at their method through arguments using batched sampling, where $N$ responses are partitioned into $N/K$ groups of size $K$. In this setting, the number of successful groups follows a binomial distribution with $M=N/K$ trials, and the arcsin transform would stabilize the variance of this estimate. Formalizing the connection between VSTs and optimization stability when applying empirical gradients to transformed rewards is open direction for future work.